# A proof of a conjecture of Erdős and Gyárfás on monochromatic path covers

### v1

@@ 1 Introduction
The vertex set of any 2-edge-coloured complete graph $K_{n}$ can be covered by two monochromatic paths.
@@
In every $2$-edge-coloured complete graph on $n$ vertices, there exists a collection of $2\sqrt{n}$ monochromatic paths, all of the same colour, which cover the entire vertex set.
@@
There exists $n_{0}\in\mathbb{N}$ such that for all $n>n_{0}$ and all 2-edge-colourings of $E(K_{n})$, there exists a collection of at most $\sqrt{n}$ monochromatic paths of the same colour that cover the vertices of $K_{n}$.
The constant $n_{0}$ that we find is large, but with some additional technical effort, one can show that for all $n\in\mathbb{N}$, $\sqrt{n}+10$ monochromatic paths of the same colour are sufficient to cover $K_{n}$.
Several related problems of this type have also been studied, looking for monochromatic structures that either cover or partition the vertex set within an edge-coloured graph. For example, Lehel’s conjecture, proven in full by Bessy and Thomassé [1], strengthens Theorem 1.1. It states that the vertices of a $2$-edge-coloured $K_{n}$ can be covered by two vertex-disjoint monochromatic cycles (of course, they must be allowed to have different colours). A famous generalisation of Lehel’s conjecture due to Gyárfás [5], stating that an $r$-edge-coloured complete graph can be covered by $r$ monochromatic cycles, remains wide open. We refer the reader to the survey [7] for partial results and further related problems.
@@ 2 Preliminaries
If $k$ and $\ell$ are distinct natural numbers then every red-blue edge-coloured complete bipartite graph $K_{\left\lceil\frac{k+l}{2}\right\rceil,\left\lceil\frac{k+l}{2}\right\rceil}$ contains a red path of length $k$ or a blue path of length $\ell$.
Suppose we have found a long monochromatic path $P$ in a red-blue edge-coloured $K_{n}$, without loss of generality we may assume $P$ is blue. We may consider the bipartite graph between $V(P)$ and a subset of $V(K_{n})\setminus V(P)$, comprised only of red edges. If there is a long red path, say $R$, is this bipartition then we obtain a large set $V(P)\cap V(R)$ covered by both a red path and a blue path, which allows us some flexibility for covering the remainder of the vertices. A similar idea may be used if we can find few red paths that cover many vertices in the bipartition. In the following two lemmas we consider certain degree properties within bipartite graphs that allow us to find few paths covering many vertices.
@@
Let $m\in\mathbb{N}$ and let $G$ be a bipartite graph on vertex set $X\cup Y$ with $|X|\geq|Y|+2m$, and such that every vertex in $Y$ has degree at least $|X|-m$. Then $G$ contains a collection of at most $\lfloor|X|/|Y|\rfloor$ paths whose union covers all but $|Y|+2m$ vertices in $X$ and all vertices in $Y$.
@@
we may apply Lemma 2.2, to obtain a path $P$ covering $2|Y|$ vertices in $G$. In particular, $P$ covers all vertices in $Y$. Let $X^{\prime}=X\setminus V(P)$, and note that in the induced subgraph $G^{\prime}=G[X^{\prime}\cup Y]$, every vertex in $Y$ has degree at least $|X^{\prime}|-m$. If $|X^{\prime}|\leq|Y|+2m$, then $\{P\}$ is a collection of paths as desired. Otherwise, we may apply the induction hypothesis to $G[X^{\prime}\cup Y]$ to obtain a collection of paths $P_{1},\ldots,P_{k}$ with $k\leq\lfloor|X^{\prime}|/|Y|\rfloor=\lfloor|X|/|Y|\rfloor-1$ that cover all but $|Y|+2m$ vertices in $X^{\prime}$. Adding $P$ to this collection yields the claim. ∎
@@
Let $G$ be a non-empty bipartite graph on vertex set $X\cup Y$ with sets $X_{0}=\{x\in X:d(x)=|Y|\}$, $X_{1}=X\setminus X_{0}$, $Y_{0}=\{y\in Y:d(y)=|X|\}$, and $Y_{1}=Y\setminus Y_{0}$ such that
@@
We prove the statement again by induction on $|X|$. It follows from (i) and (ii) that $|X_{0}|\geq|Y_{1}|+1$, and therefore, $G$ contains a path $P$ with $12|Y_{1}|+$ 1 vertices, alternating between $X_{0}$ and $Y_{1}$, such that $P$ starts at $x_{0}\in X_{0}$ and also ends in $X_{0}$. In particular $P$ covers all vertices in $Y_{1}$.
@@
Furthermore, since every vertex in $Y_{0}$ has neighbourhood $X$, and $|X|>|Y|$, $G$ contains a path $Q$ with $2\min\{|Y_{0}|,|X_{1}|\}$ vertices, alternating between $Y_{0}$ and $X\setminus V(P)$, that starts at $y_{0}\in Y_{0}$ and covers $\min\{|Y_{0}|,|X_{1}|\}$ vertices in $X_{1}$. Consider the concatenated path $R=Px_{0}y_{0}Q$ and the graph $G^{\prime}=G[X^{\prime}\cup Y]$ where $X^{\prime}=X\setminus V(R)$.
@@
Defining the sets $X_{0}^{\prime},X_{1}^{\prime},Y_{0}^{\prime}$, and $Y_{1}^{\prime}$ analogously for $G^{\prime}$ to $X_{0},X_{1},Y_{0}$, and $Y_{1}$ were defined for $G$ respectively, we observe that $Y_{1}^{\prime}\subset Y_{1}$ and $X_{1}^{\prime}\subset X_{1}\setminus V(Q)$, while $Y_{0}\subset Y_{0}^{\prime}$. If we assume that $X_{1}^{\prime}=\emptyset$ or $Y_{1}^{\prime}=\emptyset$, then it follows that $G^{\prime}$ is complete bipartite and it is trivially possible to cover $X\setminus V(R)$ with $\left\lceil\frac{|X^{\prime}|}{|Y|+1}\right\rceil=\left\lceil\frac{|X|}{|Y|+1}\right\rceil-1$ paths so that nothing is left to show. Suppose therefore that $X_{1},Y_{1}\neq\emptyset$, in particular, we must have $|X_{1}|>|Y_{0}|$ so that $|V(Q)\cap X_{1}|=|Y_{0}|$. Using this, we see that
$\frac{|X_{0}^{\prime}|}{|Y_{1}^{\prime}|}\geq\frac{|X_{0}\setminus V(P)|}{|Y_{1}|}\geq\frac{|X_{0}|}{|Y_{1}|}-2>2\left(\frac{|X_{1}|}{|Y_{0}|}-1\right)=\frac{2|X_{1}\setminus V(Q)|}{|Y_{0}|}>\frac{2|X_{1}^{\prime}|}{|Y_{0}^{\prime}|}.$
@@
If $|X^{\prime}|>|Y^{\prime}|$, we may use the induction hypothesis to obtain a family $\mathcal{P}$ of at most $\left\lceil\frac{|X|}{|Y|+1}\right\rceil-1$ paths which cover all of $X^{\prime}$ and $Y$, and together with $R$, these paths cover all of $G$.
@@
If $|X^{\prime}|\leq|Y^{\prime}|$ on the other hand, then we must have $|Y_{0}^{\prime}|>|X_{1}^{\prime}|$ so that it is possible to cover all of $X_{1}^{\prime}$ with a single path $Q^{\prime}$ of length $2|X_{1}^{\prime}|$. The remaining graph $G^{\prime}-V(Q^{\prime})$ is complete bipartite so that we may find a path $P^{\prime}$ that covers all of $X_{0}^{\prime}$ in it. Concatenating $P^{\prime}$ and $Q^{\prime}$ suitably, we obtain a second path $R^{\prime}$ such that $R$ and $R^{\prime}$ cover all of $G$. ∎
@@
Let $n\in\mathbb{N}$. In what follows, without loss of generality we assume the vertex set of $K_{n}$ is given by $[n]$. Let $\mathcal{X}_{n}$ be the set of all red-blue colourings of $E(K_{n})$. For $\chi\in\mathcal{X}_{n}$, denote by $f(n,\chi)$ the smallest family of monochromatic paths of the same colour that cover $[n]$. Furthermore, let $f(n)=\max_{\chi\in\mathcal{X}_{n}}f(n,\chi)$.
@@
Suppose there exist red paths $P_{1},\ldots,P_{k}\subset K_{n}$, blue paths $Q_{1},\ldots,Q_{k}\subset K_{n}$, and a set $S\subset[n]$ satisfying $S\subset(V(P_{1})\cup\ldots\cup V(P_{k}))\cap(V(Q_{1})\cup\ldots\cup V(Q_{k}))$. Let $H$ be the 2-edge-coloured graph obtained by deleting $S$ from $K_{n}$ under the colouring $\chi$. Then $H$ has $m=n-|S|$ vertices, and by definition of $f(m)$, there exists a collection of at most $f(m)$ monochromatic paths of the same colour which cover $V(H)$. Adding $\{P_{1},\ldots,P_{k}\}$ or $\{Q_{1},\ldots,Q_{k}\}$ to this collection according to whether these paths are red or blue respectively yields a collection of at most $f(m)+k$ monochromatic paths of the same colour covering $V(H)\cup S=[n]$ in $K_{n}$. By our assumptions, we have
@@
$\sqrt{n-|S|}\leq\sqrt{n-2(C_{1}-C_{2}+1)}\leq\sqrt{n}-(C_{1}-C_{2}+1).$
@@
Let $C_{1}\geq C_{2}\geq 0$ and let $n>8^{4}(C_{1}-C_{2}+1)^{4}$ be such that $f(m)<\sqrt{m}+C_{1}$ for all $m<n$. If $\chi\in\mathcal{X}_{n}$ is such that $f(n,\chi)>\sqrt{n}+C_{2}$, then there a monochromatic path $P$ such that the size of $Y=[n]\setminus V(P)$ is at most $\sqrt{n}+8(C_{1}-C_{2}+1)\sqrt[4]{n}$. Furthermore, if $\gamma$ is the colour of the edges of $P$, then for every $y\in Y$, the size of $\{x\in V(P):\chi(xy)=\gamma\}$ is at most $2(C_{1}-C_{2}+1)\sqrt{n}$.
@@
Let $\Delta=C_{1}-C_{2}$ and let $\chi\in\mathcal{X}_{n}$ be such that $f(n,\chi)>\sqrt{n}+C_{2}$. Without loss of generality, there exists a blue path $Q$ which covers $\lfloor n/2\rfloor$ vertices. Let $W$ be a subset of $[n]\setminus V(Q)$ of size $\lfloor n/2\rfloor$ and consider the longest red path $R$ in the complete bipartite graph between $V(Q)$ and $W$. The set $V(R)\cap V(Q)$ is covered both by a single red and a single blue path so that by Lemma 3.1, we may assume that $|V(R)\cap V(Q)|<2(\Delta+1)\sqrt{n}$. However, since $R$ is a path in a bipartite graph between $V(Q)$ and $W$, at least $\lfloor|V(R)|/2\rfloor$ vertices of $R$ must lie in $V(Q)$, and thus, it follows that $|V(R)|<4(\Delta+1)\sqrt{n}$. By Lemma 2.1, there exists a blue path of length at least $n-2(\Delta+1)\sqrt{n}-1$ between $V(Q)$ and $W$.
Let $P$ be the longest blue path in $K_{n}$, denote its vertex set by $X$, and let $Y=[n]\setminus X$. Consider $y\in Y$ and let $B$ denote the set of blue neighbours of $y$ in $P$. Since $P$ cannot be extended, $B$ does not contain any endpoint or any two consecutive vertices of $P$. Furthermore, if we direct $P$ arbitrarily, the predecessors of $B$ form a red clique set. Indeed, if $P=v_{1}v_{2}\ldots v_{k}$, $1\leq i<j\leq k$, and $v_{i}v_{j}$, $v_{i+1}y$, and $v_{j+1}y$ were all blue, then $v_{1}\ldots v_{i}v_{j}v_{j-1}\ldots v_{i+1}yv_{j+1}\ldots v_{k}$ would be a blue path of length $k+1$. However, we may argue as before that $V(P)$ cannot contain a red path of length $2(\Delta+1)\sqrt{n}$, which means that $y$ has at most $2(\Delta+1)\sqrt{n}$ blue neighbours in $X$.
In other words, taking $m=2(\Delta+1)\sqrt{n}$, each vertex in $Y$ has at least $|X|-m$ blue neighbours in $X$. What is more, we know that $|X|\geq n-2(\Delta+1)\sqrt{n}-1$ and hence $|Y|\leq 2(\Delta+1)\sqrt{n}+1$. Thus, as $n>100(\Delta+1)^{2}$,
$|X|\geq n-2(\Delta+1)\sqrt{n}-1>10(\Delta+1)\sqrt{n}-2(\Delta+1)\sqrt{n}-1>6(\Delta+1)\sqrt{n}+1\geq|Y|+2m.$
It now follows from Lemma 2.3 that the bipartite graph given by the red edges between $X$ and $Y$ contains a collection of at most $\lfloor|X|/|Y|\rfloor<n/|Y|$ paths which cover all but $|Y|+2m<6(\Delta+1)\sqrt{n}+1$ vertices in $X$ and all vertices of $Y$.
@@
$\sqrt{n}+C_{2}<\sqrt{6(\Delta+1)\sqrt{n}+1}+C_{1}+n/|Y|\leq 3(\Delta+1)\sqrt[4]{n}+C_{1}+n/|Y|,$
which we can simplify to
$\sqrt{n}<n/|Y|+4(\Delta+1)\sqrt[4]{n},$
which implies that
$|Y|<\frac{n}{\sqrt{n}-4(\Delta+1)\sqrt[4]{n}}=\frac{\sqrt{n}}{1-4(\Delta+1)/\sqrt[4]{n}}.$
Since we have $4]{n}\geq 8(\Delta+1)$ by assumption and $1/(1-x)<1+2x$ for $x\in[0,1/2]$, we get that
$|Y|<\sqrt{n}+8(\Delta+1)\sqrt[4]{n},$
@@
We prove the claim by induction on $n$, noting that it is trivial for $n\leq C$. Suppose therefore that $n>C$ and that $f(m)<\sqrt{m}+C$ for all $m<n$ and consider a colouring $\chi\in\mathcal{X}_{n}$. Applying Lemma 3.2 with $C=C_{1}=C_{2}$, we obtain without loss of generality a blue path $P$ such that the size of $Y=[n]\setminus V(P)$ is at most $\sqrt{n}+8\sqrt[4]{n}<3\sqrt{n}/2$ and such that every $y\in Y$ has at least $|X|-2\sqrt{n}$ blue neighbours in $X=V(P)$. By Lemma 3.3, if $|Y_{0}|\leq\sqrt{n}/2$ or $|Y|\leq\sqrt{n}$, then $f(n,\chi)\leq\sqrt{n}+2$ and there is nothing left to show. Let us therefore assume that $|Y_{0}|>\sqrt{n}/2$ and $|Y|>\sqrt{n}$.

### v2

@@ 1 Introduction
The vertex set of every 2-edge-coloured complete graph on $n$ vertices can be covered by two monochromatic paths.
@@
The vertex set of every 2-edge-coloured complete graph on $n$ vertices can be covered by $2\sqrt{n}$ monochromatic paths, all of the same colour.
We remark that although Theorem 1.1 still holds if we additionally impose that the two paths are vertex-disjoint, this is not the case for Theorem 1.2, where paths are allowed to intersect.
@@
For all $n>20^{40}$, the vertex set of every 2-edge-coloured complete graph on $n$ vertices can be covered by $\sqrt{n}$ monochromatic paths, all of the same colour.
We do not attempt to optimise the constant $20^{40}$, but with some additional technical effort, one can show that for all $n\in\mathbb{N}$, $\sqrt{n}+10$ monochromatic paths of the same colour are sufficient to cover $V(K_{n})$.
Several related problems of this type have also been studied, looking for monochromatic structures that either cover or partition the vertex set within an edge-coloured graph. For example, Lehel’s conjecture, proven for large $n$ by Łuczak, Rödl and Szemerédi [9] and independently by Allen [1], and for all $n$ by Bessy and Thomassé [2], strengthens Theorem 1.1. It states that the vertex set of every $2$-edge-coloured $K_{n}$ can be covered by two vertex-disjoint monochromatic cycles (of course, they must be allowed to have different colours). A generalisation of Theorem 1.1 conjectured by Gyárfás [6], stating that the vertex set of every $r$-edge-coloured complete graph can be covered by $r$ vertex-disjoint monochromatic paths, remains open. We refer the reader to the survey [8] for partial results and further related problems.
@@ 2 Preliminaries
If $k$ and $\ell$ are distinct natural numbers then every red-blue edge-coloured complete bipartite graph $K_{\left\lceil\frac{k+\ell}{2}\right\rceil,\left\lceil\frac{k+\ell}{2}\right\rceil}$ contains a red path of length $k$ or a blue path of length $\ell$, i.e., a red path on $k+1$ vertices or a blue path on $\ell+1$ vertices.
Suppose we have found a long monochromatic path $P$ in a red-blue edge-coloured $K_{n}$, without loss of generality we may assume $P$ is blue. We may consider the bipartite graph between $V(P)$ and a subset of $V(K_{n})\setminus V(P)$, comprised only of red edges. If there is a long red path, say $R$, in this bipartition then we obtain a large set $V(P)\cap V(R)$ covered by both a red path and a blue path, which allows us some flexibility for covering the remainder of the vertices. A similar idea may be used if we can find few red paths that cover many vertices in the bipartition. In the following two lemmas we consider certain degree properties within bipartite graphs that allow us to find few paths covering many vertices.
@@
Let $m\in\mathbb{N}$ and let $G$ be a bipartite graph on vertex set $X\cup Y$ with $|X|\geq|Y|+2m$, and such that every vertex in $Y$ has degree at least |X|m|X|-m. Then $G$ contains a collection of at most $\lfloor|X|/|Y|\rfloor$ paths whose union covers all but $|Y|+2m$ vertices in $X$ and all vertices in $Y$.
@@
we may apply Lemma 2.2, to obtain a path $P$ covering $2|Y|$ vertices in $G$. In particular, $P$ covers all vertices in $Y$. Let $X^{\prime}=X\setminus V(P)$, and note that in the induced subgraph $G^{\prime}=G[X^{\prime}\cup Y]$, every vertex in $Y$ has degree at least $|X^{\prime}|-m$. For the base case of our induction, if $|X^{\prime}|\leq|Y|+2m$, then $\{P\}$ is a collection of paths as desired. Otherwise, we may apply the induction hypothesis to $G[X^{\prime}\cup Y]$ to obtain a collection of paths $P_{1},\ldots,P_{k}$ with $k\leq\lfloor|X^{\prime}|/|Y|\rfloor=\lfloor|X|/|Y|\rfloor-1$ that cover all but $|Y|+2m$ vertices in $X^{\prime}$. Adding $P$ to this collection yields the claim. ∎
@@
Let $G$ be a bipartite graph on vertex set $X\cup Y$ with sets $X_{0}=\{x\in X:d(x)=|Y|\}$, $X_{1}=X\setminus X_{0}$, $Y_{0}=\{y\in Y:d(y)=|X|\}\neq\emptyset$, and $Y_{1}=Y\setminus Y_{0}$ such that
@@
We prove the statement again by induction on $|X|$, always assuming that $|X|\geq|Y|+1$. The base case $|X|=1$ is trivial, because there exists no graph $G$ as above with $1=|X|>|Y|\geq|Y_{0}|>0$. It follows from (i) and (ii) that $|X_{0}|\geq|Y_{1}|+1$. Indeed, if $X_{1}=Y_{1}=\emptyset$ then it is clear, and so supposing otherwise that $|X_{0}|/|Y_{1}|>2|X_{1}|/|Y_{0}|$ but that $|X_{0}|\leq|Y_{1}|$, then from (i) we must have $|X_{1}|>|Y_{0}|$, which cannot all hold true at the same time. Therefore assuming $|X_{0}|\geq|Y_{1}|+1$, $G$ contains a path $P$ with $2|Y_{1}|+1$ vertices, alternating between $X_{0}$ and $Y_{1}$, such that $P$ starts at $x_{0}\in X_{0}$ and also ends in $X_{0}$. In particular $P$ covers all vertices in $Y_{1}$.
@@
Furthermore, since every vertex in $Y_{0}$ has neighbourhood $X$, and $|X|>|Y|$, $G$ contains a path $Q$ with $2|Y_{0}|$ vertices, alternating between $Y_{0}$ and $X\setminus V(P)$, that starts at $y_{0}\in Y_{0}$ and covers $\min\{|Y_{0}|,|X_{1}|\}$ vertices in $X_{1}$. This is possible since $|X\setminus V(P)|=|X|-|Y_{1}|-1\geq|Y|-|Y_{1}|=|Y_{0}|$. Since $V(P)\subseteq X_{0}\cup Y_{1}$ and $V(Q)\subseteq(X\setminus V(P))\cup Y_{0}$, we have that $P$ and $Q$ are vertex-disjoint. Thus we can consider a concatenated path $R=Px_{0}y_{0}Q$ and let $X^{\prime}=X\setminus V(R)$. Note that $|X^{\prime}|=|X|-|Y|-1$.
@@
Observe that if $|X^{\prime}|=0$, i.e., if $|X|=|Y|+1$, then the collection consisting only of the path $R$ covers all vertices in $G$, as required by the lemma. Now, let us assume $|X^{\prime}|>0$ and let $G^{\prime}=G[X^{\prime}\cup Y]$. Defining the sets $X_{0}^{\prime},X_{1}^{\prime},Y_{0}^{\prime}$, and $Y_{1}^{\prime}$ analogously for $G^{\prime}$ to how $X_{0},X_{1},Y_{0}$, and $Y_{1}$ were defined for $G$ respectively, we observe that $Y_{1}^{\prime}\subseteq Y_{1}$ and $X_{1}^{\prime}=X_{1}\setminus V(Q)$, while $Y_{0}\subseteq Y_{0}^{\prime}$. In particular, $Y_{0}^{\prime}\neq\emptyset$. If we assume $X_{1}^{\prime}=\emptyset$ or $Y_{1}^{\prime}=\emptyset$, then it follows that $G^{\prime}$ is complete bipartite and it is trivially possible to cover $X\setminus V(R)$ with $\left\lceil\frac{|X^{\prime}|}{|Y|+1}\right\rceil=\left\lceil\frac{|X|}{|Y|+1}\right\rceil-1$ paths so that nothing is left to show. Suppose therefore that $X_{1}^{\prime},Y_{1}^{\prime}\neq\emptyset$. In particular, we must have $|X_{1}|>|Y_{0}|$ so that $|V(Q)\cap X_{1}|=|Y_{0}|$ and $R$ covers all of $Y$. Using the former, we see that
$\frac{|X_{0}^{\prime}|}{|Y_{1}^{\prime}|}\geq\frac{|X_{0}\setminus V(P)|}{|Y_{1}|}\geq\frac{|X_{0}|}{|Y_{1}|}-2>2\left(\frac{|X_{1}|}{|Y_{0}|}-1\right)=\frac{2|X_{1}\setminus V(Q)|}{|Y_{0}|}\geq\frac{2|X_{1}^{\prime}|}{|Y_{0}^{\prime}|}.$ (2.1)
@@
If $|X^{\prime}|>|Y|$, we may use the induction hypothesis to obtain a family $\mathcal{P}$ of at most $\left\lceil\frac{|X^{\prime}|}{|Y|+1}\right\rceil=\left\lceil\frac{|X|}{|Y|+1}\right\rceil-1$ paths which cover all of $X^{\prime}$ and $Y$, and together with $R$, these paths cover all of $G$.
@@
If $|X^{\prime}|\leq|Y|$ on the other hand, then we must have $|Y_{0}^{\prime}|>|X_{1}^{\prime}|$. Indeed if this were not the case, by 2.1 we would have $|X_{0}^{\prime}|>2|Y_{1}^{\prime}|$ implying $|X^{\prime}|=|X_{0}^{\prime}|+|X_{1}^{\prime}|>|Y_{0}^{\prime}|+2|Y_{1}^{\prime}|\geq|Y|$, a contradiction. So, assuming $|Y_{0}^{\prime}|>|X_{1}^{\prime}|$, it is possible to cover all of $X_{1}^{\prime}$ with a single path $Q^{\prime}$ of length $2|X_{1}^{\prime}|$. The remaining graph $G^{\prime}-V(Q^{\prime})$ is complete bipartite on bipartition classes of size $|X^{\prime}|-|X_{1}^{\prime}|$ and $|Y|-|X_{1}^{\prime}|$, so using $|X^{\prime}|-|X_{1}^{\prime}|\leq|Y|-|X_{1}^{\prime}|$ we may find a path $P^{\prime}$ that covers all of $X_{0}^{\prime}$ in it. Concatenating $P^{\prime}$ and $Q^{\prime}$ suitably, we obtain a second path $R^{\prime}$ such that $R$ and $R^{\prime}$ cover all of $G$. ∎
@@
Let $n\in\mathbb{N}$. In what follows, without loss of generality we assume the vertex set of $K_{n}$ is given by $[n]$. Let $\mathcal{X}_{n}$ be the set of all red-blue colourings of $E(K_{n})$. For $\chi\in\mathcal{X}_{n}$, denote by $f(n,\chi)$ the size of a smallest family of monochromatic paths of the same colour that cover $[n]$. Furthermore, let $f(n)=\max_{\chi\in\mathcal{X}_{n}}f(n,\chi)$.
@@
Suppose there exist red paths $P_{1},\ldots,P_{k}\subset K_{n}$, blue paths $Q_{1},\ldots,Q_{k}\subset K_{n}$, and a set $S\subset[n]$ satisfying $S\subset(V(P_{1})\cup\ldots\cup V(P_{k}))\cap(V(Q_{1})\cup\ldots\cup V(Q_{k}))$. Let $H$ be the 2-edge-coloured graph obtained by deleting $S$ from $K_{n}$ under the colouring $\chi$. Then $H$ has m=n|S|m=n-|S| vertices, and by definition of $f(m)$, there exists a collection of at most $f(m)$ monochromatic paths of the same colour which cover $V(H)$. Adding $\{P_{1},\ldots,P_{k}\}$ or $\{Q_{1},\ldots,Q_{k}\}$ to this collection according to whether these paths are red or blue respectively yields a collection of at most $f(m)+k$ monochromatic paths of the same colour covering $V(H)\cup S=[n]$ in $K_{n}$. By our assumptions, we have
@@
$\sqrt{n-|S|}\leq\sqrt{n-2(C_{1}-C_{2}+1)\sqrt{n}}\leq\sqrt{n}-(C_{1}-C_{2}+1).$
@@
Let $C_{1}\geq C_{2}\geq 0$ and let $n>10^{4}(C_{1}-C_{2}+1)^{4}$ be such that $f(m)<\sqrt{m}+C_{1}$ for all $m<n$. If $\chi\in\mathcal{X}_{n}$ is such that $f(n,\chi)>\sqrt{n}+C_{2}$, then there is a monochromatic path $P$ such that the size of $Y=[n]\setminus V(P)$ is at most $\sqrt{n}+10(C_{1}-C_{2}+1)\sqrt[4]{n}$. Furthermore, if $\gamma$ is the colour of the edges of $P$, then for every $y\in Y$, the size of $\{x\in V(P):\chi(xy)=\gamma\}$ is at most $2(C_{1}-C_{2}+1)\sqrt{n}$.
@@
Let $\Delta=C_{1}-C_{2}+1$ and let $\chi\in\mathcal{X}_{n}$ be such that $f(n,\chi)>\sqrt{n}+C_{2}$. By Theorem 1.1, without loss of generality, there exists a blue path $Q$ covering $\lfloor n/2\rfloor$ vertices. Let $W$ be a subset of $n]\setminus V(Q)$ of size $\lfloor n/2\rfloor$ and let $H$ be the complete bipartite graph between $V(Q)$ and $W$ under the colouring $\chi$. Let $R$ be a longest red path in $H$. The set $V(Q)\cap V(R)$ is covered both by a single red and a single blue path so that by Lemma 3.1, we may assume $|V(Q)\cap V(R)|<2\Delta\sqrt{n}$. However, since $R$ is a path in $H$, it must alternate between $V(Q)$ and $W$ and so at least $\lfloor|V(R)|/2\rfloor$ vertices of $R$ lie in $V(Q)$. We can therefore obtain the following
$\frac{|V(R)|-1}{2}\leq\left\lfloor\frac{|V(R)|}{2}\right\rfloor\leq|V(Q)\cap V(R)|<2\Delta\sqrt{n}\leq\lceil 2\Delta\sqrt{n}\rceil.$
Rearranging, we deduce that $|V(R)|<2\lceil 2\Delta\sqrt{n}\rceil+1$. Note that $H$ is a red-blue edge-coloured $K_{\lceil(n-1)/2\rceil,\lceil(n-1)/2\rceil}$, so applying Lemma 2.1 with $k=2\lceil 2\Delta\sqrt{n}\rceil$ and $\ell=n-1-k$ tells us that $H$ contains a red path on $k+1$ vertices or a blue path on $\ell+1$ vertices. Since $R$ was chosen to be a longest red path in $H$, the former case cannot occur. Thus there exists a blue path covering at least $\ell+1=n-k=n-2\lceil 2\Delta\sqrt{n}\rceil\geq n-5\Delta\sqrt{n}$ vertices in $H$.
Now let $P$ be a longest blue path in $K_{n}$, denote its vertex set by $X$, and let $Y=[n]\setminus X$. Then by the above argument, $|X|\geq n-5\Delta\sqrt{n}$ and hence $|Y|\leq 5\Delta\sqrt{n}$. Consider $y\in Y$ and let $B$ denote the set of blue neighbours of $y$ in $P$. Since $P$ cannot be extended, $B$ does not contain any endpoint or any two consecutive vertices of $P$. Furthermore, if we direct $P$ arbitrarily, the predecessors of $B$ form a red clique set. Indeed, if $P=v_{1}v_{2}\ldots v_{r}$, and $v_{i}v_{j}$, $v_{i+1}y$, and $v_{j+1}y$ were all blue for some $1\leq i<j\leq r$, then $v_{1}\ldots v_{i}v_{j}v_{j-1}\ldots v_{i+1}yv_{j+1}\ldots v_{r}$ would be a blue path covering $r+1$ vertices, contradicting the maximality of $P$. However, we may argue as before that $V(P)$ cannot contain a red path on $2\Delta\sqrt{n}$ vertices, which means that $y$ has at most $2\Delta\sqrt{n}$ blue neighbours in $X$, as required by the “furthermore” part of the lemma.
In other words, taking $m=\lfloor 2\Delta\sqrt{n}\rfloor$, each vertex in $Y$ has at least |X|m|X|-m red neighbours in $X$. Thus, as $n>196\Delta^{2}$, $|X|\geq n-5\Delta\sqrt{n}$, and $|Y|\leq 5\Delta\sqrt{n}$,
$|X|\geq n-5\Delta\sqrt{n}>14\Delta\sqrt{n}-5\Delta\sqrt{n}=9\Delta\sqrt{n}\geq|Y|+2m.$
It now follows from Lemma 2.3 that the bipartite graph given by the red edges between $X$ and $Y$ contains a collection of at most $\lfloor|X|/|Y|\rfloor<n/|Y|$ paths which cover all but $|Y|+2m\leq 9\Delta\sqrt{n}$ vertices in $X$ and all vertices in $Y$.
@@
$\sqrt{n}+C_{2}<\sqrt{14\Delta\sqrt{n}}+C_{1}+n/|Y|\leq 4\Delta\sqrt[4]{n}+C_{1}+n/|Y|.$
It follows that $\sqrt{n}<n/|Y|+5\Delta\sqrt[4]{n},$ which implies
$|Y|<\frac{n}{\sqrt{n}-5\Delta\sqrt[4]{n}}=\frac{\sqrt{n}}{1-5\Delta/\sqrt[4]{n}}.$
Since we have $\sqrt[4]{n}\geq 10\Delta$ by assumption and $1/(1-x)\leq 1+2x$ for $x\in[0,1/2]$, we obtain
$|Y|<\sqrt{n}+10\Delta\sqrt[4]{n},$
@@
We prove the claim by induction on $n$, noting that it is trivial for $n\leq C$. Suppose therefore that $n>C$ and that $f(m)<\sqrt{m}+C$ for all $m<n$ and consider a colouring $\chi\in\mathcal{X}_{n}$. Applying Lemma 3.2 with $C=C_{1}=C_{2}$, we obtain without loss of generality a blue path $P$ such that the size of $Y=[n]\setminus V(P)$ is at most $\sqrt{n}+10\sqrt[4]{n}\leq 3\sqrt{n}/2$ and such that every $y\in Y$ has at least $|X|-2\sqrt{n}$ red neighbours in $X=V(P)$. Let $Y_{0}$ be the set of vertices in $Y$ that are not contained in a blue edge with any vertex in $V(P)$, so that by Lemma 3.3, if $|Y_{0}|\leq\sqrt{n}/2$ or $|Y|\leq\sqrt{n}$, then $f(n,\chi)\leq\sqrt{n}+2$ and there is nothing left to show. Let us therefore assume that $|Y_{0}|>\sqrt{n}/2$ and $|Y|>\sqrt{n}$.